The area of a rectangular field is 160m^2 and its length is twice its width .Find the dimensions of the feild.

Answer by Engg.N Magampon(6)

A=LxW
L=2W
160m^2= W x 2W
160m^2= 2W^2
160m^2/2 = 2W^2/2
80m^2=W^2
W=square root of 80
W=8.944m
L=2W
L=2(8.944)
L=17.89m
CHECKING:
A=LxW
160m^2=8.944m x 17.89m
160m^2=159.99 or 160m^2